The perimeter of a rectangle is 94m the length is 2m more than twice the width, Find the deminsions
Answer by unlockmath(1688)

Hello,
Let's have x be the width and 2x+2 will be the length. Now we can set up an equation as:
2(2x+2)+2x=94 Rewritten as:
4x+4+2x=94 Combine like terms:
6x+4=94 Subtract 4 and divide by 6 to get:
x=15
There we go the width is 15m
and the length is 32m
Make sense?
RJ
www.math-unlock.com